# Graph Neural Networks for Identifying Steady-State Behavior in Complex Networks

### v1

@@ Abstract
In complex systems, information propagation can be defined as diffused or delocalized, weakly localized, and strongly localized. Can a machine learning model learn the behavior of a linear dynamical system on networks? In this work, we develop a graph neural network framework for identifying the steady-state behavior of the linear dynamical system. We reveal that our model learns the different states with high accuracy. To understand the explainability of our model, we provide an analytical derivation for the forward and backward propagation of our framework. Finally, we use the real-world graphs in our model for validation.
@@ I Introduction
Figure 1: Different steady-state behavior in linear dynamical systems on complex networks. We depict the nodes in the graphs with $x-y$ coordinates. We assign the sizes of a node based on the degree of a node. The $z-axis$ portrays the amount of information ($x_{i}^{*}$) on a node in the steady state. The steady-state behavior of linear dynamics on the ER random network leads to delocalization, the Scalefree network shows weak localization, and the star network shows a strong localization.
@@ II Problem Definition
Figure 2: The architecture of the Graph Neural Networks for the regression task over the graphs. The $i^{th}$ input graph (${\bf A}^{(i)}$) and the associated node features (${\bf H}^{(i)}$) are given in matrix form to the models. The Graph Neural Network (GNN) layers of the model output updated node feature matrix (${\bf H}^{{}^{\prime}(i)}$), and the readout layer gives graph level representation as $\bm{z}^{(i)}$. Further, a fully connected layer predicts the IPR value. Finally, we apply a threshold scheme Eq. (10) to identify structures associated with different linear dynamical states.
We can quantify the (de)localization behavior of the steady-state ($\bm{x}^{*}$) or the PEV ($\bm{u}\equiv\bm{u}_{1}^{\bf A}$) using the inverse participation ratio ($y_{\bm{x}^{*}}$), which is the sum of the fourth power of the state vector entries and calculate as [20]
@@
where $x^{*}_{i}$ is the $i^{\text{th}}$ component of $\bm{x^{*}}=(x^{*}_{1},x^{*}_{2},\dots,x^{*}_{n})^{T}$ and $\sum_{i=1}^{n}{x^{*}_{i}}^{2}$ is the normalization term. A vector with component $(c,c,\ldots,c)$ is delocalized and has $y_{\bm{x}^{*}}=\frac{1}{n}$ for some positive constant $c>0$, whereas the vector with components $(1,0,\ldots,0)$ yields $y_{\bm{x}^{*}}=1$ and referred as most localized. Furthermore, we consider the networks to be simple, connected and undirected. Hence, some information can easily propagate from one node to another and we never get a steady-state vector of the form $\bm{x}^{*}=(1,0,\ldots,0)$ for a connected network, and thus the IPR value lies between $\frac{1}{n}\leq y_{\bm{x}^{*}}<1$. Therefore, the localization-delocalization behavior of the linear dynamics in the network is quantified using a real value, i.e., each graph associates an IPR value [21, 20, 8, 11]. Now, to identify the states ($\bm{x}^{*}$) belong to which category of dynamical behavior for linear dynamics, we formalize a threshold scheme for identifying IPR values ($y_{\bm{x}^{*}}$) lies in the range $[1/n,1)$. We define two thresholds, $\tau_{1}$ and $\tau_{2}$, such that $11/n\leq\tau_{1}<\tau_{2}<$ 1. An additional parameter $\epsilon$ ($\epsilon>0$) defines some flexibility around the thresholds.
Delocalized Regime ($r_{1}$). IPR values significantly below the first threshold, including an $\epsilon$-width around $\tau_{1}$:
@@
Weakly localized Regime ($r_{2}$). IPR values around and between the two thresholds, including $\epsilon$-width around $\tau_{1}$ and $\tau_{2}$:
@@
Strongly localized Regime ($r_{3}$). IPR values significantly above the second threshold, including an $\epsilon$-width around $\tau_{2}$:
@@
Figure 3: Neural message passing in Graph Neural Network model for the regression task. Each node’s features are represented with a small rectangle associated with the node. (a) represents one layer of GNN where a node aggregates and updates its feature based on the immediate neighbor and its features. (b) For the second GNN layer, a node updates its features by aggregating the messages from neighbors to neighbors and its own. (c) The third layer of GNN aggregates the messages from neighbors to neighbors to neighbors. (d) The readout layer creates a graph-level representation from all node features through the mean pooling function that holds the graph’s global information based on the neural message passing framework. (e) Finally, we pass it to a fully connected neural network for the graph’s IPR value prediction task.
@@
Here, we formulate the problem as a graph regression task to predict a target value, IPR, associated with each graph structure. For a given set of graphs $\{\mathcal{G}_{i}\}_{i=1}^{N}$ where each $\mathcal{G}_{i}=(V_{i},E_{i})$ consists of a set of nodes $V_{i}$ and a set of edges $E_{i}$ such that $n_{i}=|V_{i}|$ and $m_{i}=|E_{i}|$. Each graph $\mathcal{G}_{i}$ has an associated target value, i.e., IPR value, $y^{(i)}\in\mathbb{R}$. For each node $v\in V_{i}$ in $\mathcal{G}_{i}$, there is an associated feature vector $\bm{h}_{j}^{(i)}\in\mathbb{R}^{d}$ and $\mathbf{H}^{(i)}\in\mathbb{R}^{|V_{i}|\times d}$ be the node feature matrix where $\bm{h}_{j}^{(i)}$ are the $j^{th}$ row. We represent each $\mathcal{G}_{i}$ using its adjacency matrix ${\bf A}^{(i)}$. The objective is to learn a function $f:\mathcal{G}\to\mathbb{R}$, such that $f(\mathcal{G}_{i})\approx y^{(i)}$ for the given set of $N$ graphs.
@@ III Methodology and Results
III.1 GNN Architecture
The GNN architecture comprises three graph convolutional layers, each followed by a ReLU activation function. After that, a readout layer performs mean pooling to aggregate node features into a single graph representation. Finally, we use a fully connected layer that outputs the scalar IPR value for the regression task (Fig. 2). A brief description of the architecture is provided below.
@@ III.1 GNN Architecture
Figure 5: We use scale-free networks for training and testing. (a) We can observe very low accuracy during the test time for the GCN model. (b) However, we can observe increased accuracy using the Graph Attention Network. We train the model for $500$ epochs.
Figure 6: Prediction of Graph Attention Networks (GAT). (a) We consider ER random and scale-free networks and associated IPR values leading to the dynamical states being delocalized and weakly localized. We can observe the GAT model predicts the state’s IPR value with significant accuracy. (b) Finally, the model can also predict the IPR value during testing for real-world graph data sets. We train the model for $500$ epochs for the ER and SF networks and $200$ for the real-world dataset. We choose $\tau_{1}=0.05$, $\tau_{2}=0.2$, $\epsilon=1e-6$.
@@ III.2 Data sets preparation
For the regression task, we create the graph data sets by combining delocalized, weakly localized, and strongly localized network structures, which include star, wheel, path, cycle, and random graph models as Erdős-Rényi (ER) and the Scale-free (SF) networks (Appendix D). As predictions of the GNN model are independent of network size, we vary the size of the networks during dataset creation and store them as edge lists (${\bf A}^{(i)}$). For each network, we calculate the IPR value from the principal eigenvector (PEV) and assign it as the target value ($y^{(i)}$) to $\mathcal{G}_{i}$ in the datasets. Since we do not have predefined node features for the networks, and the GCN framework requires node features as input [25], we initialize the feature vector ($\bm{h}_{j}^{(i)}$) for each node with network-specific properties (clustering coefficient, PageRank, degree centrality, betweenness centrality and closeness centrality) to form the initial feature matrix (${\bf H}^{(0,i)}$) for the $i$th graph. However, the feature matrix could also be initialized with random binary values (i.e., 00, $1$) [25]. We pass the edge indices, feature matrices, and labels into the model for the regression task, $\{({\bf A}^{(i)},{\bf H}^{(0,i)},y^{(i)})\}_{i=1}^{N}$. We primarily use small-sized networks for training the model, while testing is conducted on networks ranging from small to large sizes, including sizes similar to and beyond those used in training. We also consider a real-world data set ("NCI1") to train the model [26]. The "NCI1" contains $N=4110$ graphs, and each node has $37$ features. We preprocess the data sets, only use the first $7$ feature, and remove all the disconnected graphs and graphs of sizes less than $10$ leading to $2796$ graphs. IPR values for disconnected graphs are trivially high, and we are only considering the connected graphs for our study. For real-world data sets, we split as $80\%$ for training and $20\%$ for testing.
@@ III.3 Training and Testing Strategy
We move further and incorporate random graph structures (ER and scale-free random networks) in the data sets. Note that the ER random graph belongs to the delocalized state, and SF belongs to both the delocalized and weakly localized state. We train the model with only the SF networks, and during the testing time, one can observe accuracy is not good (Fig. 5(a)). To resolve this, we changed the model and the parameters. We choose the Graph Attention network [28], update the loss function by considering the $\log$ value, and choose AdamW optimizer instead of Adam. We also use a dropout rate of $0.6$ and set learning rare to $51e-$ 5 in the model. The new setup leads to improvement in the results (Fig. 5(b)). Now, we consider both ER and SF networks and train the model, and during the testing time, we can observe good accuracy in predicting the IPR values (Fig. 6(a)). Finally, we use real-world data to test our model; one can see the reasonable accuracy of the real-world test data sets (Fig. 6(b)).
@@ III.4 Insights of Training process
Backward Propagation: To compute the gradients for updating the weight matrices, we apply the chain rule to propagate the error from the output layer back through the network layers. We calculate the gradient of loss with respect to the output of the linear layer as
@@
Now, we calculate the gradients for the linear layer. We know each graph $i$ contributes to the overall loss $\mathcal{L}$. Therefore, we accumulate the gradient contributions from each graph when computing the gradient of the loss with respect to the weight matrix ${\bf W}^{(\text{lin})}$ (Example 2). Thus, to get the gradient of the loss with respect to the weights ${\bf W}^{(\text{lin})}$, we apply the chain rule
@@
where $\frac{\partial\hat{y}^{(i)}}{\partial{\bf W}^{(\text{lin})}}=z^{(i)}$. Similarly, we calculate gradient with respect to $b$ and $z^{(i)}$ as
@@ IV Conclusion
We introduce a framework to predict the localized and delocalized states of complex unknown networks using Graph Neural Network (GNN). We focus on leveraging the rich information embedded in network structures and extracting relevant features for graph regression. Specifically, a GCN model is employed to predict inverse participation ratio (IPR) values for unknown networks and consequently identify localized or delocalized states. Our approach provides a machine learning alternative to the traditional principal eigenvalue analysis [20] for understanding the behavior of linear dynamical processes in real-world systems at steady state. This method offers near real-time insights into the structural properties of underlying networks. A key advantage of the proposed framework is its ability to train on small networks and generalize to larger networks, achieving an accuracy of nearly $99\%$ with test unseen networks. This makes the model scale-invariant, with the computational cost of state prediction remaining consistent regardless of network size, apart from the cost of reading the network data.
@@ Acknowledgements.
PP acknowledges the Anusandhan National Research Foundation (ANRF) grant TAR/2022/000657, Govt. of India. AR is supported by the Netherlands Organisation for Scientific Research (NWO) program.

### v2

# Predicting Steady-State Behavior in Complex Networks with Graph Neural Networks

@@ Abstract
In complex systems, information propagation can be defined as diffused or delocalized, weakly localized, and strongly localized. This study investigates the application of graph neural network models to learn the behavior of a linear dynamical system on networks. A graph convolution and attention-based neural network framework has been developed to identify the steady-state behavior of the linear dynamical system. We reveal that our trained model distinguishes the different states with high accuracy. Furthermore, we have evaluated model performance with real-world data. In addition, to understand the explainability of our model, we provide an analytical derivation for the forward and backward propagation of our framework.
@@ I Introduction
Figure 1: Steady-state behavior in linear dynamical systems on complex networks. We depict the nodes in the graphs with $x-y$ coordinates. We assign the sizes of a node based on the degree of a node. The $z-axis$ portrays the amount of information ($x_{i}^{*}$) on a node in the steady state. The steady-state behavior of linear dynamics on the ER random network leads to delocalization, the Scalefree network shows weak localization, and the star network shows a strong localization.
@@ II Problem Definition
Figure 2: The architecture of the Graph Neural Networks for the regression task over the graphs. The $i^{th}$ input graph (${\bf A}^{(i)}$) and the associated node features (${\bf H}^{(i)}$) are given in matrix form to the models. The Graph Neural Network (GNN) layers of the model output updated node feature matrix (${\bf H}^{{}^{\prime}(i)}$), and the readout layer gives graph level representation as $\bm{z}^{(i)}$. Further, a fully connected layer predicts the IPR value. Finally, we apply a threshold scheme (Eq. (10)) to identify different linear dynamical states.
We quantify the (de)localization behavior of the steady-state ($\bm{x}^{*}$) or the PEV ($\bm{u}\equiv\bm{u}_{1}^{\bf A}$) using the inverse participation ratio ($y_{\bm{x}^{*}}$), which is the sum of the fourth power of the state vector entries and calculate as [20]
@@
where $x^{*}_{i}$ is the $i^{\text{th}}$ component of $\bm{x^{*}}=(x^{*}_{1},x^{*}_{2},\dots,x^{*}_{n})^{T}$ and $\sum_{i=1}^{n}{x^{*}_{i}}^{2}$ is the normalization term. A vector with component $(c,c,\ldots,c)$ is delocalized and has $y_{\bm{x}^{*}}=\frac{1}{n}$ for some positive constant $c>0$, whereas the vector with components $(1,0,\ldots,0)$ yields $y_{\bm{x}^{*}}=1$ and referred as most localized. Furthermore, we consider the networks to be simple, connected and undirected. Hence, some information can easily propagate from one node to another and we never get a steady-state vector of the form $\bm{x}^{*}=(1,0,\ldots,0)$ for a connected network, and thus the IPR value lies between $\frac{1}{n}\leq y_{\bm{x}^{*}}<1$. Therefore, the localization-delocalization behavior of the linear dynamics in the network is quantified using a real value, i.e., each graph associates an IPR value [21, 20, 8, 11]. Now, to identify the states ($\bm{x}^{*}$) belong to which category of dynamical behavior for linear dynamics, we formalize a threshold scheme for identifying IPR values ($y\equiv y_{\bm{x}^{*}}$) lies in the range $[1/n,1)$. We define two thresholds, $\tau_{1}$ and $\tau_{2}$, such that $11/n\leq\tau_{1}<\tau_{2}<$ 1. An additional parameter $\epsilon$ ($\epsilon>0$) defines some flexibility around the thresholds.
Delocalized region ($r_{1}$). IPR values significantly below the first threshold, including an $\epsilon$-width around $\tau_{1}$:
@@
Weakly localized region ($r_{2}$). IPR values around and between the two thresholds, including $\epsilon$-width around $\tau_{1}$ and $\tau_{2}$:
@@
Strongly localized region ($r_{3}$). IPR values significantly above the second threshold, including an $\epsilon$-width around $\tau_{2}$:
@@
Figure 3: Neural message passing in Graph Convolutional Networks (GCN) model for the regression task. Each node’s features are represented with a small rectangle associated with the node. (a) represents one layer of GCN where a node aggregates and updates its feature based on the immediate neighbor and its features. (b) For the second GCN layer, a node updates its features by aggregating the messages from neighbors to neighbors and its own. (c) The third layer of GCN aggregates the messages from neighbors to neighbors to neighbors. (d) The readout layer creates a graph-level representation from all node features through the mean pooling function that holds the graph’s global information based on the neural message passing framework. (e) Finally, we pass it to a fully connected neural network for the graph’s IPR value prediction task.
@@
Here, we formulate the problem as a graph regression task to predict a target value, IPR, associated with each graph structure. For a given set of graphs $\{\mathcal{G}_{i}\}_{i=1}^{N}$ where each $\mathcal{G}_{i}=(V_{i},E_{i})$ consists of a set of nodes $V_{i}$ and a set of edges $E_{i}$ such that $n_{i}=|V_{i}|$ and $m_{i}=|E_{i}|$. We represent each $\mathcal{G}_{i}$ using its adjacency matrix ${\bf A}^{(i)}$. Further, each graph $\mathcal{G}_{i}$ has an associated target value, i.e., IPR value, $y^{(i)}\in\mathbb{R}$. For each node $v\in V_{i}$ in $\mathcal{G}_{i}$, there is an associated feature vector $\bm{h}_{j}^{(i)}\in\mathbb{R}^{d}$ and $\mathbf{H}^{(i)}\in\mathbb{R}^{|V_{i}|\times d}$ be the node feature matrix where $\bm{h}_{j}^{(i)}$ is the $j^{th}$ row. The objective is to learn a function $f:\mathcal{G}\to\mathbb{R}$, such that $f(\mathcal{G}_{i})\approx y^{(i)}$ for the given set of $N$ graphs.
@@ III Methodology and Results
III.1 GCN Architecture
The GCN architecture comprises three graph convolutional layers, each followed by a ReLU activation function. After that, a readout layer performs mean pooling to aggregate node features into a single graph representation. Finally, we use a fully connected layer that outputs the scalar IPR value for the regression task (Fig. 2). A brief description of the architecture is provided below.
@@ III.1 GCN Architecture
Figure 5: We use scale-free networks for training and testing. (a) We can observe very low accuracy during the test time for the GCN model. (b) However, we can observe increased accuracy using the Graph Attention Network. (c) We consider ER random and scale-free networks and associated IPR values leading to the dynamical states being delocalized and weakly localized. We can observe the GAT model predicts the state’s IPR value with significant accuracy. Here, we train the model for $500$ epochs.
@@ III.2 Data Sets Preparation
For the regression task, we create the graph data sets by combining delocalized, weakly localized, and strongly localized network structures, which include star, wheel, path, cycle, and random graph models as Erdős-Rényi (ER) and the Scale-free (SF) networks (Appendix D). As predictions of the GNN model are independent of network size, we vary the size of the networks during dataset creation and store them as edge lists (${\bf A}^{(i)}$). For each network, we calculate the IPR value from the principal eigenvector (PEV) and assign it as the target value ($y^{(i)}$) to $\mathcal{G}_{i}$ in the datasets. Since we do not have predefined node features for the networks, and the GCN framework requires node features as input [25], we initialize the feature vector ($\bm{h}_{j}^{(i)}$) for each node with network-specific properties (clustering coefficient, PageRank, degree centrality, betweenness centrality and closeness centrality) to form the initial feature matrix (${\bf H}^{(0,i)}$) for the $i$th graph. However, the feature matrix could also be initialized with random binary values (i.e., 00, $1$) [25]. We pass the edge indices, feature matrices, and labels into the model for the regression task, $\{({\bf A}^{(i)},{\bf H}^{(0,i)},y^{(i)})\}_{i=1}^{N}$. We primarily use small-sized networks for training the model, while testing is conducted on networks ranging from small to large sizes, including sizes similar to and beyond those used in training.
[…]
We access the datasets through PyTorch Geometric libraries and preprocess the data sets by removing all the disconnected graphs and those with fewer than $10$ nodes. We incorporate node features extracted from network properties. The sizes of our preprocessed datasets are $\texttt{N}_{\texttt{ENZ}}=441$, $\texttt{N}_{\texttt{NCI}}=2796$, and $\texttt{N}_{\texttt{MCF}}=25084$ with each nodes having seven features as in the model network. Note that IPR values for disconnected graphs are trivially high, and we focus exclusively on connected graphs in our study. For real-world data sets, we divide $80\%$ of the data for training and the rest for testing.
@@ III.3 Training and Testing Strategy
We move further and incorporate random graph structures (ER and scale-free random networks) in the data sets. Note that the ER random graph belongs to the delocalized state, and SF belongs to both the delocalized and weakly localized state. We train the model with only the SF networks, and during the testing time, one can observe accuracy is not good (Fig. 5(a)). To resolve this, we changed the model and the parameters. We choose the Graph Attention network [29], update the loss function by considering the $\log$ value, and choose AdamW optimizer instead of Adam. We also use a dropout rate of $0.6$ and set learning rare to $51e-$ 5 in the model. The new setup leads to improvement in the results (Fig. 5(b)). Now, we consider both ER and SF networks and train the model, and during the testing time, we can observe good accuracy in predicting the IPR values (Fig. 5(c)).
The performance of GCN and GAT in identifying various dynamic states in model networks is highly accurate. GCN is particularly effective in distinguishing between strongly localized and delocalized states (Fig. 4), while GAT excels at differentiating weakly localized and delocalized states (Fig. 5(c)). Although trained models reliably predict states in model networks, applying them to real-world data presents challenges due to imbalanced state distribution and limited dataset size in the $r_{1}$ and $r_{3}$ regions (Fig. 6). To assess real-world applicability, we trained the GAT model on real-world data sets and achieved reasonable accuracy on test datasets (Fig. 6(a-c)).
@@ III.4 Insights of Training Process
Backward Propagation: To compute the gradients to update the weight matrices, we apply the chain rule to propagate the error from the output layer back through the network layers. We calculate the gradient of loss with respect to the output of the linear layer as
@@
Figure 7: Portray the distribution of weight matrices ($\mathbf{W}^{(0)},\mathbf{W}^{(1)},\mathbf{W}^{(2)}$) entries for the three GCN layers during the training process with cycle and star networks (Fig. 4(a)). We show the weights matrix entries for the first five epochs, where the ‘Initial Weights’ infer the initial weight matrix, and the ‘End Weights’ infer the weight matrix after the fourth epochs. The gray color indicated weight matrices during epochs $1$, $2$, and $3$.
We calculate the gradients for the linear layer. We know that each graph $i$ contributes to the overall loss $\mathcal{L}$. Therefore, we accumulate the gradient contributions from each graph when computing the gradient of the loss with respect to the weight matrix ${\bf W}^{(\text{lin})}$ (Example 2). Thus, to obtain the gradient of the loss with respect to the weights ${\bf W}^{(\text{lin})}$, we apply the chain rule
@@
where $\frac{\partial\hat{y}^{(i)}}{\partial{\bf W}^{(\text{lin})}}=z^{(i)}$. Similarly, we calculate the gradient with respect to $b$ and $z^{(i)}$ as
@@ IV Conclusion
Using the graph neural network, we introduce a framework to predict the localized and delocalized states of complex unknown networks. We focus on leveraging the rich information embedded in network structures and extracting relevant features for graph regression. Specifically, a GCN model is employed to predict inverse participation ratio values for unknown networks and consequently identify localized or delocalized states. Our approach provides a graph neural network alternative to the traditional principal eigenvalue analysis [20] for understanding the behavior of linear dynamical processes in real-world systems at steady state. This method offers near real-time insight into the structural properties of underlying networks. A key advantage of the proposed framework is its ability to train on small networks and generalize to larger networks, achieving an accuracy of nearly $\sim 100\%$ with test unseen model networks to understand delocalized or strongly localized states. This makes the model scale-invariant, with the computational cost of state prediction remaining consistent regardless of network size, apart from the cost of reading the network data.
Our trained GNN framework (e.g., GCN and GAT) effectively identifies three different states in unseen test model network data. Moreover, our model accurately identifies the states in the weakly localized regions for the real-world data. However, distinguishing between delocalized weakly localized and strongly localized states associated with real-world graphs poses a significant challenge. It might be due to the imbalance of data points in different states and limited dataset availability. Moving forward, we aim to address these challenges and improve identification accuracy.
@@ Acknowledgements.
PP is thankful to Sulthan Vishnu Sai (IIIT Raichur) for sharing the model data generation code and is indebted to Anirban Dasgupta and Shubhajit Roy (IIT Gandhinagar) for the valuable discussion. PP acknowledges the Anusandhan National Research Foundation (ANRF) grant TAR/2022/000657, Govt. of India. AR is supported by the Netherlands Organisation for Scientific Research (NWO) program.